Jerome's experiment requires him to mix a 50%-acid solution with an 80%-acid solution to create 200 mL of a 68%-acid solution. How much 50%-acid solution and how much 80%-acid solution should he use?
Answer by rfer(16322)

let x=.50 base=80ml
let 200-x=.80 base=120ml
.5x+.8(200-x)=.68*200
.5x+160-.8x=136
-.3x=-24
x=80